# A Two-Sink Solution to the Self-Similar Euler Equations

### v1

@@ 3.3 Asymptotics as $P\to 0^{-}$ of the Periods $T_{\pm}$
We finish the proof of Theorem 3.4 by showing that T=T(P)T=T(-P) decreases as P-P increases near $P=0$. The following uses the theory presented in [3].
@@
$T=\begin{cases}\displaystyle\pi\underbrace{-\frac{4\lambda^{2\lambda-2}\sqrt{\pi}\Gamma(\lambda)\tan(\pi\lambda)}{\Gamma(\lambda-\frac{1}{2})}}_{<\,0}|P|+o(|P|)&\qquad\text{if}~\lambda\in(1,\frac{3}{2}),\\ \\ \displaystyle\pi-\frac{9}{4}|P|\log(|P|^{-1})+O(|P|)&\qquad\text{if}~\lambda=\frac{3}{2},\\ \\ \displaystyle\pi\underbrace{-\frac{2^{1+\frac{1}{2\lambda-2}}\lambda\Gamma(\frac{2\lambda-1}{2\lambda-2})\Gamma(\frac{\lambda}{2\lambda-2})}{\sqrt{\pi}}\bigg(1-\sin\Big(\frac{\pi}{2\lambda-2}\Big.)\bigg)}_{<\,0}|P|^{\frac{1}{2\lambda-2}}+O(|P|)&\qquad\text{if}~\lambda\in(\frac{3}{2},2).\end{cases}$
In any cases, $T$ is a decreasing function of P-P near $P=0$.
@@ 4 Relation to the Homogeneous Shear Flow
The asymptotic behavior of the two-sink solution as $\lambda\to 2$ is of interest as a means of comparison with the power-law shear flow $\Omega(x,y)=|y|^{-\alpha}$. In particular, it appears at first sight that the two-sink approaches the power-law shear flow as $\lambda\to 2$ (equivalently $\alpha\to 0$).
@@ 4.1 Convergence to the Shear Flow as Pressure Goes to Zero
Let $\psi_{P}$ be a positive solution to the ordinary differential equation in Equation (9) with pressure $P<0$ on the interval $0,T_{+}(P)]$ corresponding to the stream function of a two-sink solution. Let $\psi_{S}(x)=\lambda^{-\lambda}\sin^{\lambda}(x)$ be the explicit shear flow solution to Equation (9) on the interval $0,\pi]$ (with zero pressure). Then we have
@@
in $C^{1,1-1/\lambda}$ on any compact subset of $0,\pi)$.
@@
However, if $\psi_{P}$ is Lipschitz, then $2P+\psi_{P}^{2-\frac{2}{\lambda}}-\lambda^{2}\psi_{P}^{2}$ is $C^{0,2-2/\lambda}$ Hölder regular, with Hölder norm depending only on $L(A)$, the uniform Lipschitz bound of $\psi_{P}$. In addition, $\pm\sqrt{-2P+\psi_{P}^{2-\frac{2}{\lambda}}-\lambda^{2}\psi_{P}^{2}}$ is $C^{0,1-1/\lambda}$ Hölder regular, with Hölder norm depending only on $L(A)$, the uniform Lipschitz bound of $\psi_{P}$. We conclude that $\psi_{P}\in C^{1,1-1/\lambda}$, with Hölder norm depending solely on $A$, the uniform bound in $P$ on $\|\psi_{P}\|_{L^{\infty}}$.
By (a straightforward modification of) the Arzelà-Ascoli theorem, we conclude that there exists a sequence $P_{m}\to 0^{-}$ such that $\psi_{m}:=\psi_{P_{m}}$ converges in $C^{1,1-1/\lambda}$ to some $C^{1,1-1/\lambda}$ function $P_{\infty}$ on any compact subset of $0,\pi]$. By the uniform boundedness, the limit is unique, in the sense that taking any sequence $P\to 0^{-}$ makes $\psi_{P}$ converge to $\psi_{\infty}$. We also remark that since $\psi_{P}\geq 0$ for each $P$, we certainly have $\psi_{\infty}\geq 0$ on $0,\pi$]. It remains to show that $\psi_{\infty}=\psi_{S}$.
@@
$(f^{\prime})^{2}+f^{2}=\lambda^{-2},$
so it is clear that $f(x)=\lambda^{-1}\sin(\lambda x)$. We conclude that $\psi_{\infty}=\psi_{S}$. ∎

### v2

@@ 3.3 Asymptotics as $P\to 0^{-}$ of the Periods $T_{\pm}$
We finish the proof of Theorem 3.3 by showing that
  • For $1<\lambda<2$, $T=T_{+}+T_{-}$ decreases as P-P increases near $P=0$, and
  • For $1<\lambda\leq\frac{3}{2}$ and $k\geq 0$, $T_{+}+kT_{-}$ decreases as P-P increases near $P=0$.
The following uses the theory presented in [3].
@@
$T=\begin{cases}\displaystyle\pi\underbrace{-\frac{4\lambda^{2\lambda-2}\sqrt{\pi}\Gamma(\lambda)\tan(\pi\lambda)}{\Gamma(\lambda-\frac{1}{2})}}_{<\,0}|P|+o(|P|)&\quad\text{if}~\lambda\in(1,\frac{3}{2}),\\ \\ \displaystyle\pi-\frac{9}{4}|P|\log(|P|^{-1})+O(|P|)&\quad\text{if}~\lambda=\frac{3}{2},\\ \\ \displaystyle\pi\underbrace{-\frac{2^{1+\frac{1}{2\lambda-2}}\lambda\Gamma(\frac{2\lambda-1}{2\lambda-2})\Gamma(\frac{\lambda}{2\lambda-2})}{\sqrt{\pi}}\bigg(1-\sin\Big(\frac{\pi}{2\lambda-2}\Big.)\bigg)}_{<\,0}|P|^{\frac{1}{2\lambda-2}}+O(|P|)&\quad\text{if}~\lambda\in(\frac{3}{2},2).\end{cases}$
In any case, $T=T_{+}+T_{-}$ is a decreasing function of P-P near $P=0$. Furthermore, in the case $1<\lambda\leq\frac{3}{2}$, as P-P increases near 0, $T_{+}$ decreases at a rate of
$\begin{cases}|P|&~\text{if}~1<\lambda<\frac{3}{2},\\ |P|\log(|P|^{-1})&~\text{if}~\lambda=\frac{3}{2},\end{cases}$
while $T_{-}$ increases at a rate of $|P|^{\frac{1}{2\lambda-2}}$. Therefore, for $1<\lambda\leq\frac{3}{2}$ and any positive integer $k$, the sum $T_{+}+kT_{-}$ decreases as P-P increases near 0.
@@ 4 Relation to the Homogeneous Shear Flow
In this section, we focus on the two-sink solution $\psi(\theta)$, which is obtained by gluing two copies each of $\psi_{+}$ and $\psi_{-}$ in the sequence $\psi_{+}$, $\psi_{-}$, $\psi_{+}$, $\psi_{-}$. The asymptotic behavior of the two-sink solution $r^{\lambda}\psi(\theta)$ as $\lambda\to 2$ is of interest as a means of comparison with the power-law shear flow $\Omega(x,y)=|y|^{-\alpha}$ (where $\alpha=2-\lambda$). In particular, it appears at first sight that the two-sink approaches the power-law shear flow as $\lambda\to 2$ (equivalently $\alpha\to 0$).
@@ 4.1 Convergence to the Shear Flow as Pressure Goes to Zero
Let $\psi_{P}$ be a positive solution to the ordinary differential equation in Equation (8) with pressure $P<0$ on the interval $0,T_{+}(P)]$ corresponding to the stream function of a two-sink solution. Let $\psi_{S}(x)=\lambda^{-\lambda}\sin^{\lambda}(\theta)$ be the explicit shear flow solution to Equation (8) on the interval $0,\pi]$ (with zero pressure). Then we have
@@
in $C^{1,\gamma}$ on any compact subset of $0,\pi)$ and any $\gamma<1-1/\lambda$.
@@
Since $\psi_{P}$ is Lipschitz, then $2P+\psi_{P}^{2-\frac{2}{\lambda}}-\lambda^{2}\psi_{P}^{2}$ is $C^{0,2-2/\lambda}$ Hölder regular, with Hölder norm depending only on $L(A)$, the uniform Lipschitz bound of $\psi_{P}$. In addition, $\pm\sqrt{-2P+\psi_{P}^{2-\frac{2}{\lambda}}-\lambda^{2}\psi_{P}^{2}}$ is $C^{0,1-1/\lambda}$ Hölder regular, with Hölder norm depending only on $L(A)$, the uniform Lipschitz bound of $\psi_{P}$. We conclude that $\psi_{P}\in C^{1,1-1/\lambda}$, with Hölder norm depending solely on $A$, the uniform bound in $P$ on $\|\psi_{P}\|_{L^{\infty}}$.
By (a straightforward modification of) the Arzelà-Ascoli theorem, we conclude that there exists a sequence $P_{m}\to 0^{-}$ such that $\psi_{m}:=\psi_{P_{m}}$ converges in $C^{1,\gamma}$ to some $C^{1,1-1/\lambda}$ function $\psi_{\infty}$ on any compact subset of $0,\pi]$. If we prove that the limit is unique, taking any sequence $P\to 0^{-}$ makes $\psi_{P}$ converge to $\psi_{\infty}$. We also remark that since $\psi_{P}\geq 0$ for each $P$, we certainly have $\psi_{\infty}\geq 0$ on $0,\pi$]. It remains to show that $\psi_{\infty}=\psi_{S}$.
@@
$(f^{\prime})^{2}+f^{2}=\lambda^{-2}.$
Since $f(0)=0$ and $f\geq 0$, it is clear that $f(x)=\lambda^{-1}\sin(\theta)$. We conclude that $\psi_{\infty}=\psi_{S}$. ∎